OpenStudy (pphalke):

is (1,-6) a solution to the equation 5p-7q=47? need help

Vocaloid (vocaloid):

hint: if p = 1 and q = -6 what is 5p-7q?

OpenStudy (pphalke):

-37 so its not a solution

Vocaloid (vocaloid):

check your math again

Vocaloid (vocaloid):

5-(-42) = ?

OpenStudy (pphalke):

what is -7(-6)?

Vocaloid (vocaloid):

5p-7q = 5(1) - 7(-6) = 5 - (-42) = ?

Vocaloid (vocaloid):

you forgot the other negative sign in your calculations. you need to include the negative sign on 6 AND the negative sign in front of 7